Samson & Me is a very popular hair salon. It offers high-quality hair-styling and physical relaxation services at a reasonable price, so it always has unlimited demand. The service process includes these activities that are conducted in the sequence described below. (The time required for each activity is shown in parenthesis):

  • Activity 1: Welcome a guest and offer homemade herb tea. (8 minutes)
  • Activity 2: Wash and condition hair. (10 minutes)
  • Activity 3: Design the hairstyle and style the hair. (28 minutes)

Three servers (S1, S2, and S3) offer the services, and the assignment of tasks to servers is as follows:

  • S1 does Activity 1.
  • S2 does Activity 2.
  • S3 does Activity 3.

In the Samson & Me hair salon, suppose that Activities 1 and 3 take 5 and 18 minutes, respectively. Also suppose that S1 now does both Activity 1 and Activity 2 while either S2 or S3 can perform Activity 3. (Note that this means Activity 3 requires only one of S2 and S3, not both.) What is the capacity rate of the whole process, in units per hour?
Answer #1

Installed capacity of Samson & Me Hair Salon= Time taken in Activities 1, 2 and 3 = 8+10+28= 46 minutes per unit

Again actual capacity of Samson & Me Hair Salon = Actual time taken in Activities 1, 2 and 3 by Server S1 and S2 or S3= 5+10+18= 33 minutes per unit

Therefore capacity rate of the whole process of Hair styling and physical relaxation in units per hour = Actual capacity/Installed capacity *60 = 33/46*60= 43.04 = 43 units per hour (Omitting the decimal values).